OpenStudy (anonymous):

Hailyn has $120 in savings. She plans to add $25 to her savings each month. How much will she have in savings after 9 months? A. $370 B. $345 C. $154 D. $145

OpenStudy (anonymous):

B: is the right answer

OpenStudy (texaschic101):

m = number of months 120 + 25m = total amount after m months when m = 9 120 + 25(9) = total (T) 120 + 225 = T 345 = T yep...I agree...its B

OpenStudy (anonymous):

Thank you! :)